# Reconstructions of Einstein-Aether Gravity from Barrow Agegraphic and New Barrow Agegraphic Dark Energy models: Examinations and Observational Limits

### v1

@@ Abstract
We present a comprehensive investigation exploring the theoretical framework of Einstein-Aether gravity theory when combined with two novel cosmological paradigms: the Barrow Agegraphic Dark Energy (BADE) and its newer variant, the New Barrow Agegraphic Dark Energy (NBADE). Our study focuses on deriving the functional relationships within Einstein-Aether gravity as they emerge from these dark energy formulations. The parameter space of our theoretical models is rigorously constrained through statistical analysis employing the Markov Chain Monte Carlo (MCMC) methodology, utilizing multiple observational datasets, incorporating measurements from cosmic chronometers (CC), Baryon Acoustic Oscillations (BAO), and the combined Pantheon+SH0ES compilation. Based on our optimized parameter sets, we conduct an extensive analysis of fundamental cosmological indicators, including cosmographic parameter evolution, dark energy equation of state parameter ($\omega_{DE}$), evolution of the density parameter $\Omega(z)$, dynamical characteristics in the $\omega^{\prime}_{DE}-\omega_{DE}$ space, behavior of statefinder diagnostic pairs $(r,s^{*})$ and $(r,q)$, and Om(z) diagnostic trajectories. Our analysis demonstrates that the current cosmic expansion exhibits accelerated behavior, with the dark energy component manifesting quintessence-like properties in the present epoch while trending toward phantom behavior in future evolution. We additionally evaluate the viability of both BADE and NBADE frameworks through an examination of the squared sound speed ($v_{s}^{2}$) stability criterion. The cumulative evidence suggests that these models effectively characterize contemporary cosmic evolution while offering novel perspectives on dark energy phenomenology.
@@ 1 Introduction
A different approach to addressing the issue of cosmic acceleration is by altering the geometric component of the Einstein-Hilbert action. This modification can potentially account for the observed accelerated expansion without the necessity of introducing an ad hoc dark energy component (For detailed reviews, refer to [26, 27, 28]). The idea of altering the standard theory of gravity was first investigated through $f(R)$ gravity, which extends the Einstein-Hilbert action by incorporating a function of the Ricci scalar curvature [29]. This pioneering work has inspired the development of numerous modified gravity models, including but not limited to those discussed in [30, 31, 32, 33, 34, 35, 36]. For a more extensive review of modified gravitational models, refer to [37, 38, 26]. Einstein-Aether gravity is another intriguing modified gravity model, which introduces a unit, time-like vector field, referred to as the “aether”, alongside the tensor field in Einstein’s equations [39]. Several interesting works on Einstein-Aether gravity are described in [40, 41, 42, 43, 44].
@@
In the field of cosmology, the conventional approach involves solving field equations to derive the space-time geometry. However, the reconstruction method presents an alternative paradigm wherein the underlying geometry is presumed based on observational evidence or theoretical considerations, and subsequently, model equations consistent with empirical observations are derived [45, 46, 47, 48, 49]. Recent advancements in theoretical physics have sparked a growing interest in investigating possible connections between different Holographic Dark Energy (HDE) models and various modified gravity theories. This field of research has garnered significant attention, as it presents promising prospects for shedding light on the mysterious nature of dark energy and its far-reaching implications for cosmology. The intersection of HDE models and modified gravity theories offers a rich area for exploration, with the capability to offer a deeper insight into the fundamental structure of the universe and the processes responsible for its accelerated expansion. By examining the interplay between these theoretical frameworks, researchers aim to construct more comprehensive and robust models that can reconcile observational data with theoretical predictions, thereby elevating our concept of the evolution of the universe and the fundamental aspects of dark energy. The evolution of Einstein-Aether gravity research has seen significant developments through various theoretical frameworks. A pioneering study was conducted examining both conventional and entropy-corrected versions of holographic and new agegraphic dark energy models within the Einstein-Aether gravity context [50]. This investigation provided crucial insights into the reconstructed function’s behavior and its cosmological significance. Further theoretical advancement came through the analysis of Einstein-Aether gravity’s reconstruction from diverse modified gravity theories, with particular attention to their cosmological stability characteristics [51]. Subsequent research established meaningful connections between Einstein-Aether gravity and various scalar field dark energy models, revealing important temporal patterns in both scalar fields and their associated potentials throughout cosmic history [52]. The theoretical framework was further expanded through investigations incorporating multiple holographic dark energy variants, including Tsallis, Renyi, and Sharma-Mittal formulations [53]. These studies utilized various analytical tools, such as the equation of state parameter $\omega_{DE}$ and the $\omega^{\prime}_{DE}-\omega_{DE}$ phase plane, to explore their cosmological implications. The theoretical foundation for these holographic approaches originates from ’t Hooft’s groundbreaking work [54], which built upon fundamental principles of black hole thermodynamics [55, 56]. This holographic principle proposes a fundamental relationship between volumetric space information and its boundary representation, conceptually similar to how two-dimensional holograms capture three-dimensional information. Many researchers now consider the Holographic principle fundamental in the quantum description of gravity. Applications of Holographic dark energy models have been explored extensively [57, 58, 59, 60, 61, 62]. Recently, significant attention has been drawn to the newly introduced Barrow entropy correction for black holes [63]. Barrow Holographic Dark Energy (BHDE), as developed by Saridakis [64], emerges from this corrected entropy and offers promising improvements over conventional holographic dark energy models [65, 66, 67]. The motivation for this study lies in the broader applicability of Barrow entropy, which provides a more generalized framework than the Bekenstein-Hawking entropy. Barrow entropy’s resemblance to Tsallis non-extensive entropy [68, 69, 70] further underscores its relevance, while in a specific limit, it recovers the standard Bekenstein-Hawking entropy. This quantum-gravitational deformation correction to the black hole horizon may offer a more accurate description of cosmological phenomena, potentially resolving long-standing puzzles in the field. As demonstrated in Ref. [71], BHDE effectively captures the thermal evolution of the universe throughout the periods dominated by dark energy and dark matter. Moreover, recent work has shown that Barrow Entropic Dark Energy is a subset of a more general Holographic Dark Energy framework when an appropriate IR cutoff is chosen [72]. A recent modification inspired by Barrow entropy was proposed in [73], where the infrared (IR) cutoff is selected as the universe’s age and the conformal time, leading to the development of Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) models. The authors have shown several interesting cosmological features in their model. The confluence of these theoretical and observational developments motivates our investigation into the reconstruction of the functional forms of Einstein-Aether gravity using BADE and NBADE as the background evolution. This study employs a specific formulation of the characteristic length to analyze the resulting cosmological implications. Our research methodology incorporates observational data from multiple sources to limit the parameters of our proposed model:
@@
By leveraging this comprehensive empirical dataset, we aim to establish robust observational bounds on the model parameters, thereby facilitating a more profound understanding of the underlying physical scenarios. To achieve this, we employ the covariance matrix technique, which enables us to minimize associated errors and determine optimal values for several key parameters within our proposed model. This multi-faceted approach, combining CC, BAO, and Pantheon+SH0ES data, provides a more comprehensive and statistically significant basis for our analysis. The inclusion of the Pantheon+SH0ES dataset, with its extensive 1701 data points, significantly enhances the statistical power of our study, allowing for more precise constraints on the model parameters. Our methodology not only aims to reconstruct the functional forms of BADE and NBADE but also seeks to elucidate their cosmological implications as per the recent and comprehensive observational data available. The organization of this paper is as follows: in Sect. 2 We have reviewed some important cosmological parameters in Sect. 3 we have presented the Einstein-Aether gravity in Sect. 4 we have discussed in brief about the Barrow agegraphic dark energy in Sect. 4.1 and Sect. 4.2 we have reconstructed Einstein-Aether gravity from BADE and NBADE. Sect. 5 We have discussed the methodology used in our analysis, Sect. 6 is dedicated to the findings of our analysis, and finally, we end our paper with concluding remarks in Sect. 7.
@@ 4 Barrow Agegraphic Dark Energy
A novel approach to dark energy modeling was introduced by Sharma, Varshney, and Dubey [73], who formulated two distinct frameworks: the Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) models. These frameworks distinguish themselves through their respective infrared (IR) cutoff selections, with BADE utilizing the universe’s age $T$ and NBADE employing the conformal time $\eta$. These parameters are mathematically expressed as:
@@
In these expressions, $T$ represents the universal age parameter, while $\eta$ denotes the conformal time measure. Our investigation proceeds under the assumption of non-interacting dark matter and dark energy components.
@@ 4.1 Reconstruction from BADE
Here, we will derive the unknown function $F(K)$ from Barrow Agegraphic Dark Energy (BADE). Now, using the relations in (33) and (38) we obtained,
@@ 4.2 Reconstruction from NBADE
Now, we will reconstruct the Einstein-Aether gravity from New Barrow Agegraphic Dark Energy (NBADE). From the relations given in (39), (33) and taking the above defined scale factor relation $a(t)=a_{0}t^{\delta}$, the conformal time $\eta$ can be written as,
@@
Here, $C_{2}$ is the integration constant. The equation (49) gives the reconstructed expression of $F(K)$ from New Barrow Agegraphic Dark Energy. The above function $F(K)$ is analytic except $M=0$, $K=0$, $\delta=1$ and $\Delta=\dfrac{1-2\delta}{\delta-1}$. Thus reconstruction of Einstein-Aether gravity from NBADE is feasible as finite number of singularities are there. We have shown the plot of the reconstructed $F(K)$ against $K$ in the fig. 1(b) for several values of $\Delta$. As it is seen from the plot, the function $F(K)$ is decreasing with the increase of $K$ when it is reconstructed from New Barrow Agegraphic Dark Energy.

### v2

@@ Abstract
We present a comprehensive investigation exploring the theoretical framework of Einstein-Aether gravity theory when combined with two modified cosmological paradigms: the Barrow Agegraphic Dark Energy (BADE) and its newer variant, the New Barrow Agegraphic Dark Energy (NBADE). Our study focuses on reconstructing the functional form of the Einstein-Aether Lagrangian component $F(K)$ from these phenomenological dark energy models. Model parameters are constrained using a Markov Chain Monte Carlo (MCMC) approach based on multiple datasets, including cosmic chronometers (CC), Baryon Acoustic Oscillations (BAO), and the Pantheon+SH0ES compilation. Using best-fit parameters, we analyze various cosmological diagnostics: Hubble and deceleration parameter evolution, dark energy equation of state ($\omega_{DE}$), density parameter trajectories, $\omega_{DE}^{\prime}$$\omega_{DE}$ phase space behavior, statefinder diagnostics $(r,s^{*})$ and $(r,q)$, and Om(z) trajectories. Both models exhibit late-time acceleration, with the dark energy sector showing a quintessence-like nature in the current epoch and evolving toward a phantom regime in the future. Stability analysis based on the squared sound speed ($v_{s}^{2}$) highlights partial epoch-dependent stability. While our results demonstrate reasonable agreement with observational data and reveal physically plausible dynamics, the models do not yet offer a fundamentally superior alternative to other dark energy reconstructions. Nonetheless, their behavior under modified entropy assumptions and their flexibility in dynamical diagnostics provide a useful framework for probing non-standard extensions of Einstein-Aether gravity and dark energy phenomenology.
@@ 1 Introduction
A different approach to addressing the issue of cosmic acceleration is by altering the geometric component of the Einstein-Hilbert action. This modification can potentially account for the observed accelerated expansion without the necessity of introducing an ad hoc dark energy component (For detailed reviews, refer to [26, 27, 28]). The idea of altering the standard theory of gravity was first investigated through $f(R)$ gravity, which extends the Einstein-Hilbert action by incorporating a function of the Ricci scalar curvature [29]. This pioneering work has inspired the development of numerous modified gravity models, including but not limited to those discussed in [30, 31, 32, 33, 34, 35, 36]. For a more extensive review of modified gravitational models, refer to [37, 38, 26]. Einstein-Aether gravity represents a theoretically significant modified gravity model that introduces a unit, time-like vector field (the “aether”) alongside the metric tensor in Einstein’s equations [39]. This approach is fundamentally motivated by considering potential Lorentz symmetry violations at high energies, which various quantum gravity approaches suggest might exist. The aether field establishes a preferred reference frame for the universe, effectively reintroducing an absolute time concept while preserving general covariance of the theory’s mathematical structure. This characteristic makes Einstein-Aether theory particularly relevant for cosmological applications, as it naturally aligns with the cosmic time coordinate implicit in the FLRW metric. Several significant developments in Einstein-Aether gravity have emerged through studies investigating black hole solutions [40], cosmological perturbations [41], axionic extensions [42], and diverse cosmological applications [43, 44]. Of particular relevance to our current study, Einstein-Aether theory provides a natural framework for accommodating time-dependent cosmological phenomena, making it an ideal candidate for integration with time-based dark energy models such as those derived from Barrow entropy considerations.
@@
In the field of cosmology, the conventional approach involves solving field equations to derive the space-time geometry. However, the reconstruction method presents an alternative paradigm wherein the underlying geometry is presumed based on observational evidence or theoretical considerations, and subsequently, model equations consistent with empirical observations are derived [45, 46, 47, 48, 49]. Recent advancements in theoretical physics have sparked a growing interest in investigating possible connections between different Holographic Dark Energy (HDE) models and various modified gravity theories [50, 51, 52, 53]. This field of research has garnered significant attention, as it presents promising prospects for shedding light on the mysterious nature of dark energy and its far-reaching implications for cosmology. The intersection of HDE models and modified gravity theories offers a rich area for exploration, with the capability to offer a deeper insight into the fundamental structure of the universe and the processes responsible for its accelerated expansion. By examining the interplay between these theoretical frameworks, researchers aim to construct more comprehensive and robust models that can reconcile observational data with theoretical predictions, thereby elevating our concept of the evolution of the universe and the fundamental aspects of dark energy. The evolution of Einstein-Aether gravity research has seen significant developments through various theoretical frameworks. A pioneering study was conducted examining both conventional and entropy-corrected versions of holographic and new agegraphic dark energy models within the Einstein-Aether gravity context [54]. This investigation provided crucial insights into the reconstructed function’s behavior and its cosmological significance. Further theoretical advancement came through the analysis of Einstein-Aether gravity’s reconstruction from diverse modified gravity theories, with particular attention to their cosmological stability characteristics [55]. Subsequent research established meaningful connections between Einstein-Aether gravity and various scalar field dark energy models, revealing important temporal patterns in both scalar fields and their associated potentials throughout cosmic history [56]. The theoretical framework was further expanded through investigations incorporating multiple holographic dark energy variants, including Tsallis, Renyi, and Sharma-Mittal formulations [57]. These studies utilized various analytical tools, such as the equation of state parameter $\omega_{DE}$ and the $\omega^{\prime}_{DE}-\omega_{DE}$ phase plane, to explore their cosmological implications. The theoretical foundation for these holographic approaches originates from ’t Hooft’s groundbreaking work [58], which built upon fundamental principles of black hole thermodynamics [59, 60]. This holographic principle proposes a fundamental relationship between volumetric space information and its boundary representation, conceptually similar to how two-dimensional holograms capture three-dimensional information. Many researchers now consider the Holographic principle fundamental in the quantum description of gravity. Applications of Holographic dark energy models have been explored extensively [61, 62, 63, 64, 65, 66]. Recently, significant attention has been drawn to the newly introduced Barrow entropy correction for black holes [67]. Barrow Holographic Dark Energy (BHDE), as developed by Saridakis [68], emerges from this corrected entropy and offers promising improvements over conventional holographic dark energy models [69, 70, 71]. The motivation for this study lies in the broader applicability of Barrow entropy, which provides a more generalized framework than the Bekenstein-Hawking entropy. Barrow entropy’s resemblance to Tsallis non-extensive entropy [72, 73, 74] further underscores its relevance, while in a specific limit, it recovers the standard Bekenstein-Hawking entropy. This quantum-gravitational deformation correction to the black hole horizon may offer a more accurate description of cosmological phenomena, potentially resolving long-standing puzzles in the field. As demonstrated in Ref. [75], BHDE effectively captures the thermal evolution of the universe throughout the periods dominated by dark energy and dark matter. Moreover, recent work has shown that Barrow Entropic Dark Energy is a subset of a more general Holographic Dark Energy framework when an appropriate IR cutoff is chosen [76]. A recent modification inspired by Barrow entropy was proposed in [77], where the infrared (IR) cutoff is selected as the universe’s age and the conformal time, leading to the development of Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) models. The authors have shown several interesting cosmological features in their model. A particularly important aspect of our study is the comparative analysis between BADE and NBADE models. While both incorporate Barrow entropy corrections, they differ fundamentally in their choice of infrared cutoff: BADE employs the universe’s age ($T$), while NBADE utilizes conformal time ($\eta$). This distinction is not merely a mathematical convenience but represents different physical interpretations of the relevant cosmological scale. The conformal time cutoff in NBADE offers several theoretical advantages over the conventional age parameter: it tracks the maximum comoving distance that light could have traveled since the Big Bang, thus providing a more complete accounting of the universe’s causal structure. This choice also addresses a critical issue in the original agegraphic models—namely that models using cosmic age ($T$) as the IR cutoff encounter difficulties in reproducing the matter-dominated era [78]. Furthermore, conformal time naturally connects to the particle horizon, which has stronger theoretical foundations in quantum gravitational approaches to cosmology. Our comparative reconstruction of Einstein-Aether gravity from both models allows us to evaluate which IR cutoff yields more physically consistent results when confronted with observational data, potentially revealing deeper insights into the nature of quantum gravity effects on cosmological scales. The synergy between Einstein-Aether gravity and Barrow Agegraphic Dark Energy models presents a particularly compelling theoretical framework for several reasons. Einstein-Aether theory’s introduction of a preferred cosmic time direction through its timelike vector field naturally complements the time-based infrared cutoffs fundamental to Agegraphic dark energy models. This temporal alignment provides a consistent theoretical foundation for investigating cosmic evolution across different epochs. Moreover, both frameworks address different aspects of the same fundamental cosmological puzzles: while Einstein-Aether gravity modifies the geometric sector to account for potential Lorentz invariance violations at high energies, Barrow entropy introduces quantum gravitational corrections to conventional holographic principles. By reconstructing Einstein-Aether gravity within the context of Barrow Agegraphic models, we can explore how quantum-gravitational deformations of entropy might manifest within a modified gravitational framework that already accommodates a preferred cosmic time. This approach not only offers new perspectives on the accelerated expansion problem but also provides potential insights into the quantum nature of spacetime itself, possibly bridging phenomenological cosmology with more fundamental physical principles. Previous studies have explored various holographic dark energy models within Einstein-Aether framework [54, 57], but the unique quantum-gravitational aspects of Barrow entropy and its implications for Einstein-Aether reconstruction remain unexplored territory with significant theoretical promise. The confluence of these theoretical and observational developments motivates our investigation into the reconstruction of the functional forms of Einstein-Aether gravity using BADE and NBADE as the background evolution. The selection of BADE and NBADE as frameworks for reconstructing Einstein-Aether gravity is motivated by several theoretical considerations. Unlike conventional dark energy models, these formulations incorporate quantum gravitational effects through Barrow’s entropy corrections, which account for potential fractal structures in black hole horizons. This quantum-gravitational foundation aligns conceptually with the motivations behind Einstein-Aether theory, which similarly seeks to address potential fundamental modifications to spacetime structure. Furthermore, the age of the universe and conformal time, which serve as infrared cutoffs in BADE and NBADE respectively, provide natural cosmological scales that can be meaningfully interpreted within Einstein-Aether’s preferred reference frame. The time-based nature of these cutoffs makes them particularly suitable for reconstruction within a framework that already incorporates a preferred cosmic time direction. This theoretical congruence suggests that exploring the reconstruction of Einstein-Aether gravity from these specific dark energy models may yield insights that would be inaccessible through other combinations of modified gravity and dark energy frameworks. Our research methodology incorporates observational data from multiple sources to limit the parameters of our proposed model:
@@
By leveraging this comprehensive empirical dataset, we aim to establish robust observational bounds on the model parameters, thereby facilitating a more profound understanding of the underlying physical scenarios. The novelty of our approach lies in the unique combination of Einstein-Aether gravity with Barrow entropy-based agegraphic models, creating a theoretical framework that simultaneously addresses potential Lorentz violations and quantum gravitational effects in cosmological evolution. To achieve this, we employ the covariance matrix technique, which enables us to minimize associated errors and determine optimal values for several key parameters within our proposed model. This multi-faceted approach, combining CC, BAO, and Pantheon+SH0ES data, provides a more comprehensive and statistically significant basis for our analysis than previous studies in this domain. The inclusion of the Pantheon+SH0ES dataset, with its extensive 1701 data points, significantly enhances the statistical power of our study, allowing for more precise constraints on the model parameters and potentially revealing new insights into the interplay between modified gravity and quantum-gravitationally motivated dark energy. Our methodology not only aims to reconstruct the functional forms of Einstein-Aether gravity from BADE and NBADE but also seeks to elucidate how these frameworks might collectively address fundamental questions about the nature of spacetime, cosmic acceleration, and the quantum structure of the universe. The organization of this paper is as follows: in Sect. 2 We have reviewed some important cosmological parameters in Sect. 3 we have presented the Einstein-Aether gravity in Sect. 4 we have discussed in brief about the Barrow agegraphic dark energy in Sect. 4.1 and Sect. 4.2 we have reconstructed Einstein-Aether gravity from BADE and NBADE. Sect. 5 We have discussed the methodology used in our analysis, Sect. 6 is dedicated to the findings of our analysis, and finally, we end our paper with concluding remarks in Sect. 7.
@@ 4 Barrow Agegraphic Dark Energy
[…]
The combined framework of conformal time as an IR cutoff and Barrow entropy corrections provides a more coherent theoretical picture when viewed through the lens of emergent gravity theories [111]. In these approaches, gravity emerges as an entropic force related to information holographically encoded on surfaces. The conformal time naturally connects to the lightcone structure of spacetime, which is fundamental to defining causal relationships and information bounds in quantum gravitational frameworks. Moreover, recent developments in quantum gravity approaches such as causal set theory [114] and causal dynamical triangulations [115] have emphasized the importance of causal structure in quantum spacetime, which aligns more naturally with the conformal time parameter than with proper time. The conformal time captures the maximum potential causal influence throughout cosmic history, making it a more fundamental quantity from an information-theoretic perspective. When these quantum gravitational considerations are combined with Einstein-Aether theory’s introduction of a preferred timelike direction, the NBADE framework offers a unique opportunity to probe potential signatures of quantum gravity in cosmological observations. The reconstruction of Einstein-Aether gravity from NBADE thus represents not merely an exercise in fitting cosmological data but a deeper exploration of how fundamental quantum gravitational principles might manifest in the large-scale structure and evolution of the universe. The cosmic age and conformal time are mathematically expressed as:
@@
In these expressions, $T$ represents the universal age parameter, while $\eta$ denotes the conformal time measure. To provide a clearer theoretical motivation for studying both BADE and NBADE within the Einstein-Aether framework, a comparative analysis of their key characteristics is given in Tab. 4:
Table 4: Comparative analysis of BADE and NBADE characteristics
Aspect BADE (Age Cutoff) NBADE (Conformal Time Cutoff)
Causal Structure Limited to proper time measurements Captures full light-cone causal structure
Matter Era Compatibility Challenges in reproducing matter-dominated epoch Successfully accommodates matter-dominated phase
Theoretical Consistency May require additional mechanisms for early universe Naturally connects early and late-time dynamics
Quantum Gravity Alignment Limited connection to quantum causal structure Better aligned with quantum spacetime frameworks
Einstein-Aether Compatibility Aligns with preferred time direction Provides richer connection to spacetime causal structure
This comparison reveals that NBADE is not merely an ad hoc variant of BADE but represents a theoretically motivated refinement that addresses specific limitations of the original formulation. The conformal time cutoff in NBADE offers a more comprehensive framework for understanding dark energy dynamics across all cosmic epochs, while maintaining the quantum-gravitational insights provided by Barrow entropy. Like ADE[116], depends on the cosmic age $T$, which varies significantly across different epochs, making the model strongly epoch-dependent and less suitable as a universal framework. In the early universe, it behaves like a cosmological constant but does not dominate; during the matter-dominated era, its scaling behavior creates inconsistencies, preventing it from naturally overtaking matter. In contrast, NBADE functions effectively as a single-parameter model, with all physical quantities determined by the parameter $n$, offering analytical simplicity and greater predictive power[117]. Furthermore, when $n$ is of order unity, NBADE inherently addresses the cosmological coincidence problem, providing a more consistent and observationally viable alternative[118]. This theoretical advantage makes the reconstruction of Einstein-Aether gravity from NBADE particularly compelling, as it combines three conceptually aligned elements: modified gravity with a preferred time direction (Einstein-Aether), quantum-gravitational deformations of entropy (Barrow entropy), and a cosmologically complete causal structure (conformal time cutoff). The comparison between reconstructions from BADE and NBADE within the same Einstein-Aether framework allows us to evaluate which IR cutoff better captures the underlying physics of dark energy when both observational constraints and theoretical consistency are considered. This approach provides a robust methodology for discriminating between competing dark energy models based on both empirical evidence and theoretical foundations. Our investigation proceeds under the assumption of non-interacting dark matter and dark energy components.
@@ 4.1 Reconstruction from BADE
Here, we will derive the unknown function $F(K)$ from BADE. Now, using the relations in (33) and (38) we obtained,
@@ 4.2 Reconstruction from NBADE
Now, we will reconstruct the Einstein-Aether gravity from NBADE. From the relations given in (39), (33) and taking the above defined scale factor relation $a(t)=a_{0}t^{\delta}$, the conformal time $\eta$ can be written as,
@@
Here, $C_{2}$ is the integration constant. The equation (49) gives the reconstructed expression of $F(K)$ from NBADE. The above function $F(K)$ is analytic except $M=0$, $K=0$, $\delta=1$ and $\Delta=\dfrac{1-2\delta}{\delta-1}$. Thus reconstruction of Einstein-Aether gravity from NBADE is feasible as finite number of singularities are there. We have shown the plot of the reconstructed $F(K)$ against $K$ in the fig. 1(b) for several values of $\Delta$. As it is seen from the plot, the function $F(K)$ is decreasing with the increase of $K$ when it is reconstructed from New Barrow Agegraphic Dark Energy.